# Game-Theoretic Joint Incentive and Cut Layer Selection Mechanism in Split Federated Learning

### v1

@@ VII Conclusion and Future Work
This paper has studied the interaction between an SFL model owner and clients during the SFL process, formulated as a single-leader multi-follower Stackelberg game in competitive or conflict situations. Each client’s strategy involves the amount of data contributed for local training, while the SFL model owner determines suitable incentives to motivate client participation in SFL. The SFL model owner also sets the cut layer to balance the training burden between clients and servers while trying to meet the required privacy level. Our results indicate that Stackelberg strategies are desirable operating points for all participants in competitive scenarios. For a more practical approach, future work involves designing a more sophisticated incentive management game that considers unshared information between clients, detailed evaluations of client contributions, and computing and networking dynamics (such as a Stackelberg Bayesian game-based and deep reinforcement learning-based incentive management).

### v3

@@ VII Conclusion and Future Work
This paper has studied the interaction between an SFL model owner and clients during the SFL process, formulated as a single-leader multi-follower Stackelberg game in competitive or conflict situations. Each client’s strategy involves the amount of data contributed for local training, while the SFL model owner determines suitable incentives to motivate client participation in SFL. The SFL model owner also sets the cut layer to balance the training burden between clients and servers while trying to meet the required privacy level. Our results indicate that Stackelberg strategies are desirable operating points for all participants in competitive scenarios. For a more practical approach, future work may involve designing a more sophisticated incentive management game that considers unshared information between clients, detailed evaluations of client contributions, and computing and networking dynamics (such as a Stackelberg Bayesian game-based or deep reinforcement learning-based incentive management).